OpenStudy (anonymous):

pls help ... For very small angles of projection 'α' with horizontal , the ratio of the range to the maximum height of a particle projected upwards from level ground with initial speed 'u' is a. 4/α b. 4u/gα

OpenStudy (anonymous):

For small angles, sinα = α using the range equation \[range = u ^{2}\sin(2\alpha)/g = u ^{2}2\alpha/g\] The velocity at max height is 0, so\[height = (u _{f}^{2} - u _{i}^{2})/(-2g) = (usin( \alpha))^{2}/2g = u ^{2} \alpha ^{2}/2g\] so the ratio is\[(u ^{2}2\alpha/g)/(u ^{2} \alpha ^{2}/2g) = 4/\alpha\]

OpenStudy (anonymous):

tnx a lot...i'm sorry i couldnt thank u earlier...i thought this question was not answered..